OpenStudy (anonymous):

Select the possible locations for the graph of (x, y), in terms of axes, given the condition xy = 0 a. x- axis c. origin b. y- axis d. all of these

OpenStudy (anonymous):

the answer is all of these. is xy = 0, then one of them have to be 0. but, there also can be multiple answers: like x = 5 and y = 0, 5(0) = 0, but x is at 5, on the x axis. same for y, y = 5 and x = 0, (0)5 = 0, but y is on 5. if both are 0, then (0)0 = 0 so all 3 are correct.

OpenStudy (anonymous):

For the function describe the domain and range. f(x)=x^2 a.domain: all imaginary numbers range: all negative numbers c.domain: all real numbers range: all non negative numbers b.domain: all imaginary numbers range: all non-negative numbers d.domain: all real numbers range: all negative numbers

OpenStudy (anonymous):

for f(x) = x^2, the answer is D. it can answer all real numbers and all negitive numbers.

OpenStudy (anonymous):

sorry, domain: all real #'s, range: all negitive numbers

OpenStudy (anonymous):

Graph the following functions,f(x)=x^2,g(x)=1/2x^2 , and identify the transformation from f to g. a.horizontal compression c.horizontal stretch b.vertical stretch d.vertical compression @SauceBoss72

OpenStudy (anonymous):

vertical compression. g(x) graph is wide, but when you square it by putting G(x) into f(x) equation, it vertical compresses it.

OpenStudy (anonymous):

Write a direct variation equation that relates x and y. Assume that y varies directly as x. Then solve. If y = –5 when x = 25, find x when y = 45

OpenStudy (anonymous):

The length of the sides of four triangles are given. Determine which triangle is not a right triangle. a. 5 mm, 12 mm, 13 mm b. 20 mm, 48 mm, 52 mm c. 6 mm, 8 mm, 10 mm d. 11 mm, 24 mm, 26 mm

OpenStudy (anonymous):

@SauceBoss72

OpenStudy (anonymous):

for the direct variation equation is 1/5th of a difference, so x is equal to 9

OpenStudy (anonymous):

for your second question, the answer is D. all of the first 3 are right triangles if you use the pythagorean therom

OpenStudy (anonymous):

How many $5 movie tickets can you buy with $37? a. 7 c. 6 b. 8 d. 9 @SauceBoss72

OpenStudy (anonymous):

the answer is A, 7. just do 37 - 5 until you cant do it anymore. youll get 2 when you subtract 7 times, or just do 7 x 5 = 35 37-35 = 2

OpenStudy (anonymous):

Given F(x)=x^20-1 , write an ordered pair (x, y) for the following elements of the domain. x = 0 a. (1, -1) c. (0, -1) b. (0, 1) d. (0, 0)

OpenStudy (anonymous):

@SauceBoss72

OpenStudy (anonymous):

domain is (1, -1)

OpenStudy (anonymous):

Which of the following transformations have been applied to the graph of the parent function,y=x^2, to produce the graph of the function y=0.4(x+2)^2+5 a. horizontal translation of -2, vertical translation of 5, vertical compression by a factor of 0.4 b. reflection across the x-axis, horizontal translation of -2, vertical translation of 5, vertical stretch by a factor of 4 c. horizontal translation of 2, vertical translation of -5, vertical compression by a factor of 0.4 d. reflection across the x-axis, horizontal translation of -2, vertical translation of -5, vertical stretch by a factor of 4

OpenStudy (anonymous):

@SauceBoss72

OpenStudy (anonymous):

How many $5 raffle tickets do you have to sell to raise $32 a. 7 c. 6 b. 8 d. 9

OpenStudy (anonymous):

for the transformations, its A, and the raffle question, its 7.